# Université de Genève, CH-1211 Genève, Switzerland

### v1

@@ 3.1 General conjectures
${\tilde{D}}_{\pm s}(q)=\sum_{k|s}{\mu(k)\over k}a_{\pm{s\over k}}(q^{k})\,,\quad s\in{\mathbb{Z}}_{>0}\,.$ (3.49)
Note that this is the formal inverse of Eq. (3.45), which can be equivalently written as
$a_{\pm n}(q)=\sum_{k|n}{1\over k}{\tilde{D}}_{\pm{n\over k}}(q^{k})\,,\quad n\in{\mathbb{Z}}_{>0}\,.$ (3.50)
We comment that, in both examples of the figure-eight and the three-twist knots studied in Section 4, we have precise expressions for $\sigma_{\alpha,m}$ for all choices of $\alpha\neq\rm{geom}$, which are given in Eqs. (4.140), (4.160), and (4.161). Moreover, our statements above simplify notably in the particular case of the geometric branch, whose function $\phi^{(\rm{geom}),\rm{WKB}}(X;q)$ has a series expansion including only the positive powers of $X$ and whose coefficients $a^{(\rm{geom})}_{n}(q)$ have a trivial factor of $q^{\sigma_{{\rm geom},n}}=1$ in the denominator.
Let us explain how our conjectures originate from insights from the TS/ST correspondence. See mmrev for a review and further references. In the context of the TS/ST correspondence, the analog of the AJ conjecture for the wave function in Eq. (2.20) is given by the quantization of the mirror curve to a toric Calabi–Yau threefold, which defines a spectral problem. Its wave function can be computed in the WKB approximation as a formal power series in $\hbar$. It was noted in kpamir ; mz-wv that this $\hbar$-expansion can be suitably resummed, producing the same picture that we have described in our claims above. This structure is similar to the one appearing in perturbative open topological string theory ov ; lmv , but it is slightly less constrained. Specifically, one has the structure of ov but not the more detailed substructure found in lmv . The integrality property in Eq. (3.44) defines what is called an admissible series in ks  [^5] We thank Stavros Garoufalidis for pointing this out to us. and is found in other examples like, e.g., in as .
We can now ask how the resummed, perturbative partition functions in Eq. (3.39) relate to the exact partition function $\chi(u;\hbar)$ described in Section 2. It follows both from physical arguments bdp and from explicit calculations gk-qseries ; dimofte-levelk ; ggm2 that the exact wave function can be decomposed as
@@
where the sum runs over the different branches of the classical $A$-polynomial labeled by $\alpha$, $\mathsf{C}_{\alpha}$ is an appropriate complex constant, and ${\cal S}^{(\alpha)}(X;q)$ is known as holomorphic block. Since the holomorphic block must reproduce the perturbative expansion of the wave function, it can be identified with the resummed WKB series in Eq. (3.39), that is,
${\cal S}^{(\alpha)}(X;q)=\exp\left(\frac{{\rm i}}{\hbar}s^{(\alpha)}_{0}(x)+s^{(\alpha)}_{1}(x)+\phi^{(\alpha),\rm{WKB}}(X;q)\right)\,.$ (3.52)
Therefore, our conjecture in Eq. (3.44) implies that holomorphic blocks in complex CS theory have the stated integrality/multicovering property, i.e., they are admissible series in the Kontsevich–Soibelman sense. This could have been anticipated from the relation between holomorphic blocks and open topological string partition functions in some examples bdp . Indeed, we will show in Section 4 that, starting from the known closed expression for the holomorphic and antiholomorphic blocks, we can successfully verify the integrality structure of Eq. (3.44) in the examples of the figure-eight and the three-twist knots for all choices of the saddle connection $\alpha$.
Let us comment that the structure in Eqs. (3.51), (3.52) is again similar to what has been found in the open TS/ST correspondence kpamir ; mz-wv ; mz-wv2 ; szabolcs ; fgrassi : the exact wave functions are given by sums of products of the resummed WKB wave functions and their duals [^6] In the TS/ST correspondence, one can also define off-shell wave functions, in which case the dual blocks differ from the original ones.. As we mentioned in Section 1, a structure similar to the one in Eq. (3.40) has already been conjectured in ekholm , where the resummed WKB expansion is obtained by solving directly the quantized $A$-polynomial equation. The approach of ekholm does not appeal to the decomposition into holomorphic blocks, and our explicit results for the resummation of the WKB expansion in the case of, e.g., the figure-eight knot in Section 4 appear to be different from those in ekholm  [^7] One reason for this difference, which was pointed out to us by Sergei Gukov, is that the results of ekholm involve the super $A$-polynomial, and they might lead to a different specialization for $\text{SL}(2,{\mathbb{C}})$..
@@ 3.2 Cancellation of singularities
Again, we hide the explicit dependence on the classical solution $\alpha$ to simplify the notation. Here and in the rest of this work, we will only re-introduce it when necessary. Recall that the exact wave function $\chi(u;\hbar)$ is well-defined for $\hbar>0$. In particular, it is well-defined for
@@
where $P,Q$ are coprime positive integers. We will refer to values of $\hbar$ of the form in Eq. (3.53) as rational values. On the other hand, due to the conjectured integrality structure in Eq. (3.44), $\phi^{\rm WKB}(X;q)$ is singular for these values of $\hbar$ precisely, thus implying that the singularities have to cancel in the sum
$\phi^{\rm WKB}(X;q)+\phi^{\rm WKB}(X_{D};q_{D})\,,$ (3.54)
which appears in the formula for $\chi(u;\hbar)$ in Eq. (3.51). This leads to non-trivial constraints on the coefficients $\hat{a}_{\pm n}(q)$ in Eq. (3.41). This type of cancellation of singularities played a major role in the understanding of the ABJM matrix model hmo2 and the TS/ST correspondence km ; ghm ; wzh .
We will now study how this cancellation occurs in the exact wave function $\chi(u;\hbar)$ as given in Eq. (3.51), and, in Section 3.3, we will compute the finite, well-defined piece which is left from the cancellation and constitutes the sum in Eq. (3.54). Moreover, as we will see following szabolcs in Section 3.4, it is possible to compute this finite part by using only the information contained in the quantum operator $\hat{A}(\hat{P},\hat{X},q)$. We take
$q={\rm e}^{{\rm i}\hbar}={\rm e}^{2\pi{\rm i}{P\over Q}}\,,\quad P,Q\in{\mathbb{Z}}_{>0}\;\;\text{coprime}\,,$ (3.55)
and, substituting into Eq. (3.40), we find that the possible singularities of $\phi^{\rm WKB}(X;q)$ occur in $a_{\pm n}(q)$ when $n=sQ$ and $s\in{\mathbb{Z}}_{>0}$. For these values of $n$, the coefficient functions $a_{\pm n}(q)$ in Eq. (3.41) become
$a_{\pm sQ}\left(q\right)=\frac{\hat{a}_{\pm sQ}\left(q\right)}{sQ(q^{sQ}-1)q^{\sigma_{\pm sQ}}}\,.$ (3.56)
@@
$q_{D}={\rm e}^{{4\pi^{2}\over\hbar}}={\rm e}^{2\pi{\rm i}{Q\over P}}\,,$ (3.57)
the possible singularities of the dual function $\phi_{\rm WKB}(X_{D};q_{D})$ occur in $a_{\pm n}(q_{D})$ when $n=sP$ and $s\in{\mathbb{Z}}_{>0}$. Indeed, the coefficient functions $a_{\pm n}(q_{D})$ in Eq. (3.41) become
$a_{\pm sP}\left(q_{D}\right)=\frac{\hat{a}_{\pm sP}\left(q_{D}\right)}{sP(q_{D}^{sP}-1)q_{D}^{\sigma_{\pm sP}}}\,.$ (3.58)
@@
$q_{\epsilon}={\rm e}^{{\rm i}\hbar_{\epsilon}}\,,\quad q_{\epsilon,D}={\rm e}^{{4\pi^{2}{\rm i}\over\hbar_{\epsilon}}}\,,\quad\hbar_{\epsilon}=2\pi{P\over Q}+\epsilon\,,\quad 0<\epsilon\ll 1\,,$ (3.59)
and consider the limit $\epsilon\rightarrow 0$. As detailed in Appendix A, Eqs. (3.56), (3.58) produce the NLO $\epsilon$-expansions
$a_{\pm sQ}\left(q_{\epsilon}\right)$ $=\left(-\frac{{\rm i}}{(sQ)^{2}\epsilon}-\frac{1}{2sQ}+\frac{1}{(sQ)^{2}}q\frac{\partial}{\partial q}\right)\frac{\hat{a}_{\pm sQ}(q)}{q^{\sigma_{\pm sQ}}}+{\cal O}(\epsilon)\,,$ (3.60a)
$a_{\pm sP}\left(q_{\epsilon,D}\right)$ $=\left(\frac{{\rm i}}{(sQ)^{2}\epsilon}+\frac{1}{sP}\left(\frac{{\rm i}}{2\pi sQ}-{1\over 2}\right)+\frac{1}{(sP)^{2}}q_{D}\frac{\partial}{\partial q_{D}}\right)\frac{\hat{a}_{\pm sP}(q_{D})}{q_{D}^{\sigma_{\pm sP}}}+{\cal O}(\epsilon)\,.$ (3.60b)
We then require the cancellation of the $\epsilon$-poles in Eqs. (3.60a), (3.60b), which yields the relation
$\hat{a}_{\pm sQ}\left({\rm e}^{2\pi{\rm i}{P\over Q}}\right){\rm e}^{-2\pi{\rm i}{P\over Q}\sigma_{\pm sQ}}=\hat{a}_{\pm sP}\left({\rm e}^{2\pi{\rm i}{Q\over P}}\right){\rm e}^{-2\pi{\rm i}{Q\over P}\sigma_{\pm sP}}\,,$ (3.61)
for all $s,P,Q\in{\mathbb{Z}}_{>0}$ with $P,Q$ coprime. We remark that, for the geometric branch, the formula in Eq. (3.61) assumes the simplified form [^8] A similar cancellation requirement was obtained in szabolcs in the study of the wave function for quantum mirror curves.
$\hat{a}^{(\rm{geom})}_{sQ}\left({\rm e}^{2\pi{\rm i}{P\over Q}}\right)=\hat{a}^{(\rm{geom})}_{sP}\left({\rm e}^{2\pi{\rm i}{Q\over P}}\right)\,,$ (3.62)
@@
Let us now prove that the cancellation formula in Eq. (3.61) is indeed a consequence of the integrality structure presented in Eq. (3.44). More precisely, Eqs. (3.41), (3.45) imply that
$\hat{a}_{n}(q)q^{-\sigma_{n}}=\sum_{k|n}kD_{k}\left(q^{n\over k}\right)q^{-{n\over k}\sigma_{k}}\sum_{j=0}^{k-1}q^{{n\over k}j}\,,$ (3.63)
@@
$\hat{a}_{sQ}\left({\rm e}^{2\pi{\rm i}{P\over Q}}\right){\rm e}^{-2\pi{\rm i}{P\over Q}\sigma_{sQ}}=\sum_{k|sQ}kD_{k}\left({\rm e}^{2\pi{\rm i}\frac{sP}{k}}\right){\rm e}^{-2\pi{\rm i}\frac{sP}{k}\sigma_{k}}\sum_{j=0}^{k-1}{\rm e}^{2\pi{\rm i}\frac{sP}{k}j}\,.$ (3.64)
Since the sum over $j$ on the RHS is non-zero if and only if $k|s$, Eq. (3.64) gives
$\hat{a}_{sQ}\left({\rm e}^{2\pi{\rm i}{P\over Q}}\right){\rm e}^{-2\pi{\rm i}{P\over Q}\sigma_{sQ}}=\sum_{k|s}k^{2}D_{k}(1)=c_{s}\in{\mathbb{Z}}\,,\quad s\in{\mathbb{Z}}_{>0}\,,$ (3.65)
which is independent of the choice of $P,Q\in{\mathbb{Z}}_{>0}$ coprime, thus proving the cancellation identity in Eq. (3.61). We note that the case of $\hat{a}_{-n}(q)$ is handled analogously, yielding the same result. In particular, recalling that $\sigma_{-n}=0$, $n\in{\mathbb{Z}}_{>0}$, it defines the sequence of integers
$\hat{a}_{-sQ}\left({\rm e}^{2\pi{\rm i}{P\over Q}}\right)=\sum_{k|s}k^{2}D_{-k}(1)=c_{-s}\in{\mathbb{Z}}\,,\quad s\in{\mathbb{Z}}_{>0}\,.$ (3.66)
@@
$D_{\pm s}(1)={1\over s^{2}}\sum_{k|s}\mu(k)c_{\pm{s\over k}}\,,\quad s\in{\mathbb{Z}}_{>0}\,,$ (3.67)
@@ 3.3 The wave function in the rational case, I
We will now apply the results obtained in Section 3.2 to derive explicitly the exact wave function $\chi(u;\hbar)$ in Eq. (3.51) as a power series in $X,X^{-1}$ for rational values of $\hbar$. Let us start from the holomorphic components $\phi^{(\alpha),{\rm WKB}}(X;q)$ and fix $\hbar$ as in Eq. (3.53), that is,
$q={\rm e}^{{\rm i}\hbar}={\rm e}^{2\pi{\rm i}{P\over Q}}\,,\quad P,Q\in{\mathbb{Z}}_{>0}\;\;\text{coprime}\,.$ (3.68)
As before, we drop the label $\alpha$ for simplicity and we intend the following discussion to hold for each choice of the classical branch independently. The singular terms come from the values $n=sQ$ for $s\in{\mathbb{Z}}_{>0}$, while choosing $n=sQ+k$ with $11\leq k\leq Q-$ 1 produces only regular terms. Specifically, the terms involving singularities are obtained by substituting Eq. (3.60a) into Eq. (3.40) and contain the non-singular contributions
$\sum_{s\geq 1}\left(-{1\over 2}+{1\over sQ}q{\partial\over\partial q}\right)\frac{\hat{a}_{\pm sQ}(q)}{sQq^{\sigma_{\pm sQ}}}X^{\pm sQ}\,.$ (3.69)
@@
$\sum_{k=1}^{Q-1}\frac{1}{q^{k}-1}\sum_{s\geq 0}{\hat{a}_{\pm(sQ+k)}(q)\over(sQ+k)q^{\sigma_{\pm(sQ+k)}}}X^{\pm(sQ+k)}\,,$ (3.70)
where we have used that $q^{sQ+k}=q^{k}$. Thus, the sum of the quantities in Eqs. (3.69), (3.70) is the finite part of $\phi^{\rm WKB}(X;q)$ evaluated at the rational point in Eq. (3.53). It is convenient to introduce the functions
@@
$\varphi^{(+)}_{k}(X)$ $=\sum_{s\geq 0}{\hat{a}_{sQ+k}(q)\over(sQ+k)q^{\sigma_{sQ+k}}}X^{sQ+k}\,,\quad k=1,\dots,Q-1\,,$ (3.71c)
$\varphi^{(-)}_{k}(X)$ $=\sum_{s\geq 0}{\hat{a}_{-(sQ+k)}(q)\over-(sQ+k)q^{sQ+k+\sigma_{-(sQ+k)}}}X^{-(sQ+k)}\,,\quad k=1,\dots,Q-1\,,$ (3.71d)
@@
$\varphi_{k}(X)$ $=\varphi^{(+)}_{k}(X)-q^{k}\varphi^{(-)}_{k}(X)\,,\quad k=1,\dots,Q-1\,.$ (3.72b)
Putting Eqs. (3.69), (3.70) together and suitably identifying the functions just introduced in Eq. (3.3), we can write the holomorphic block as
@@
Let us now consider the antiholomorphic component $\phi^{\rm WKB}(X_{D};q_{D})$ and fix $\hbar$ as above. The derivation in this case is very similar to the previous one, but the novelty is the appearance of $\log X$ terms, which originate from the $\epsilon$-expansion of the dual variable
@@
Let us add a small positive term $\epsilon$ to the rational value of $\hbar$ as in Eq. (3.59). Then, we consider the limit $\epsilon\rightarrow 0$ and find that
$X_{\epsilon,D}=X^{2\pi\over\hbar_{\epsilon}}=X^{{Q\over P}{1\over 1+{\epsilon\over 2\pi}{Q\over P}}}=X^{Q\over P}\left(1-{Q^{2}\over 2\pi P^{2}}\epsilon\log X+{\cal O}(\epsilon^{2})\right)\,.$ (3.75)
As described before, the singular terms come from the values $n=sP$ for $s\in{\mathbb{Z}}_{>0}$, while choosing $n=sP+k$ with $11\leq k\leq P-$ 1 produces only regular terms. Specifically, the terms involving singularities are obtained by substituting Eqs. (3.60b), (3.75) into the Eq. (3.40) and contain the non-singular contributions
$\sum_{s\geq 1}\left(\frac{{\rm i}}{2\pi sQ}-{1\over 2}+{1\over sP}q_{D}{\partial\over\partial q_{D}}\mp\frac{{\rm i}}{2\pi}\log X\right)\frac{\hat{a}_{\pm sP}(q_{D})}{sPq_{D}^{\sigma_{\pm sP}}}X^{\pm sQ}\,.$ (3.76)
@@
$\sum_{k=1}^{P-1}\frac{1}{q_{D}^{k}-1}\sum_{s\geq 0}{\hat{a}_{\pm(sP+k)}(q_{D})\over(sP+k)q_{D}^{\sigma_{\pm(sP+k)}}}X^{\pm(sP+k){Q\over P}}\,,$ (3.77)
where we have used that $q_{D}={\rm e}^{2\pi{Q\over P}}$ and thus $q_{D}^{sP+k}=q_{D}^{k}$. Therefore, the sum of the quantities in Eqs. (3.76), (3.77) is the finite part of the function $\phi^{\rm WKB}(X_{D};q_{D})$ evaluated at the rational point in Eq. (3.53). As for the holomorphic block, we introduce the functions

### v2

@@ 3.1 General conjectures
${\tilde{D}}_{\pm s}(q)=\sum_{k|s}{\mu(k)\over k}a_{\pm{s\over k}}\bigl(q^{k}\bigr),\qquad s\in{\mathbb{Z}}_{>0}.$ (3.11)
Note that this is the formal inverse of equation (3.7), which can be equivalently written as
$a_{\pm n}(q)=\sum_{k|n}{1\over k}{\tilde{D}}_{\pm{n\over k}}\bigl(q^{k}\bigr),\qquad n\in{\mathbb{Z}}_{>0}.$ (3.12)
Let us explain how our conjectures originate from insights from the TS/ST correspondence. See [37] for a review and further references. In the context of the TS/ST correspondence, an analogous statement to the AJ conjecture for the CS wave function in equation (2.21) arises from the quantization of the mirror curve to a toric Calabi–Yau threefold and the corresponding spectral problem. Its wave function can be computed in the WKB approximation as a formal power series in $\hbar$. It was noted in [31, 38] that this $\hbar$-expansion could be suitably resummed, producing the same picture we described in our claims above. This structure is similar to the one appearing in perturbative open topological string theory [36, 40], but it is slightly less constrained. Specifically, one has the structure of [40] but not the more detailed substructure found in [36]. The integrality property in equation (3.6) characterizes what is called an admissible series by Kontsevich and Soibelman in [34], [^7] We thank Stavros Garoufalidis for pointing this out to us. and it is observed in other instances, e.g., as shown in [1].
We can now ask how the resummed perturbative partition functions in equation (3.1) relate to the exact partition function $\chi(u;\hbar)$ described in Section 2. It follows both from physical arguments [5] and from explicit computations [8, 17, 19] that the exact wave function can be conjecturally decomposed as
@@
where the sum runs over the different non-abelian branches of the classical $A$-polynomial labeled by $\alpha$, $\mathsf{C}_{\alpha}$ is an appropriate complex constant, and ${\cal S}^{(\alpha)}(X;q)$ is known as holomorphic block. Since the holomorphic block labeled by $\alpha$ must reproduce the perturbative expansion of the wave function around the corresponding classical solution, it can be identified with the resummed WKB series in equation (3.1), that is,
${\cal S}^{(\alpha)}(X;q)=\exp\left({{{\rm i}\over\hbar}}s^{(\alpha)}_{0}(x)+s^{(\alpha)}_{1}(x)+\phi^{(\alpha),{\rm WKB}}(X;q)\right).$ (3.14)
Therefore, our conjecture in equation (3.6) implies that holomorphic blocks in complex CS theory have the stated integrality/multicovering-type property, i.e., they are admissible series in the Kontsevich–Soibelman sense. This could have been anticipated from the relation between holomorphic blocks and open topological string partition functions found in some examples [5]. Indeed, we will show in Section 4 that, starting from the known closed expressions for the holomorphic and antiholomorphic blocks [5, 8, 17], we can successfully verify the integrality structure in equation (3.6) for all non-abelian choices of the flat connection in the examples of the figure-eight and three-twist knots. Furthermore, we will provide a compact proof of admissibility in the same examples by direct application of a theorem of [34] together with well-established $q$-series identities.
Let us comment that the structure in equations (3.13) and (3.14) is again similar to what has been found in the open TS/ST correspondence [15, 31, 38, 39, 44]: the exact wave functions are given by sums of products of the resummed WKB wave functions and their duals. [^8] In the TS/ST correspondence, one can also define off-shell wave functions, in which case the dual blocks differ from the original ones. As we mentioned in Section 1, a structure similar to the one in equation (3.2) has already been conjectured in [12], where the resummed WKB expansion is obtained by solving directly the quantized $A$-polynomial equation. The approach of [12] does not appeal to the decomposition into holomorphic blocks, and our explicit results for the resummation of the WKB expansion in the case of, e.g., the figure-eight knot in Section 4 appear to be different from those in [12]. [^9] One reason for this difference, which was pointed out to us by Sergei Gukov, is that the results of [12] involve the super $A$-polynomial, which might lead to a different specialization for $\mathrm{SL}(2,{\mathbb{C}})$.
@@ 3.2 Cancellation of singularities
Recall that the exact wave function $\chi(u;\hbar)$ is a holomorphic function of $\hbar\in\mathbb{C}^{\prime}$. In particular, it is well-defined for
@@
where $P$, $Q$ are coprime positive integers. We will refer to values of $\hbar$ of the form in equation (3.15) as rational values. On the other hand, due to the conjectured integrality structure in equation (3.6), $\phi^{\rm(\alpha),WKB}(X;q)$ is singular precisely for these values of $\hbar$, thus implying that the singularities must disappear in the decomposition of $\chi(u;\hbar)$ in equation (3.13). More precisely, we will show how these singularities cancel in the sum
$\phi^{\rm(\alpha),WKB}(X;q)+\phi^{\rm(\alpha),WKB}(X_{D};q_{D}),$ (3.16)
which implies non-trivial constraints on the coefficients $\hat{a}_{\pm n}^{(\alpha)}(q)$ in equation (3.3). This type of branch-by-branch cancellation of singularities found a prototypical example in Faddeev’s quantum dilogarithm [18] and played a major role in the understanding of the ABJM matrix model [24] and the TS/ST correspondence [21, 29, 42].
Again, we hide the explicit dependence on the classical solution $\alpha$ to simplify the notation. Here and in the rest of this work, we will only reintroduce it when necessary. We will now study how the cancellation occurs in the exact wave function $\chi(u;\hbar)$ as given in equation (3.13), and, in Section 3.3, we will compute the finite, well-defined piece that is left from the cancellation and constitutes the sum in equation (3.16). Moreover, as we will see in Section 3.4 along the lines of [44], it is possible to compute this finite part by using only the information contained in the quantum operator $\hat{A}\bigl(\hat{P},\hat{X},q\bigr)$. We take
$q={\rm e}^{{\rm i}\hbar}={\rm e}^{2\pi{\rm i}{P\over Q}},\qquad P,Q\in{\mathbb{Z}}_{>0}\ {\rm coprime},$ (3.17)
and, substituting into equation (3.2), we find that the possible singularities of $\phi^{\rm WKB}(X;q)$ occur when $n=sQ$ and $s\in{\mathbb{Z}}_{>0}$. In fact, for these values of $n$, the coefficient functions $a_{\pm n}(q)$ in equation (3.3) become
$a_{\pm sQ}(q)={{\hat{a}_{\pm sQ}(q)\over sQ\bigl(q^{sQ}-1\bigr)q^{\sigma_{\pm sQ}}}}.$ (3.18)
@@
$q_{D}={\rm e}^{{4\pi^{2}{\rm i}\over\hbar}}={\rm e}^{2\pi{\rm i}{Q\over P}},$ (3.19)
the possible singularities of the dual function $\phi^{\rm WKB}(X_{D};q_{D})$ occur when $n=sP$ and $s\in{\mathbb{Z}}_{>0}$. Indeed, the coefficient functions $a_{\pm n}(q_{D})$ in equation (3.3) become
$a_{\pm sP}(q_{D})={{\hat{a}_{\pm sP}(q_{D})\over sP\bigl(q_{D}^{sP}-1\bigr)q_{D}^{\sigma_{\pm sP}}}}.$ (3.20)
@@
$q_{\epsilon}={\rm e}^{{\rm i}\hbar_{\epsilon}},\qquad q_{\epsilon,D}={\rm e}^{{4\pi^{2}{\rm i}\over\hbar_{\epsilon}}},\qquad\hbar_{\epsilon}=2\pi{P\over Q}+\epsilon,\qquad 0<\epsilon\ll 1,$ (3.21)
and consider the limit $\epsilon\rightarrow 0$. As detailed in Appendix A, equations (3.18) and (3.20) produce the NLO $\epsilon$-expansions
$a_{\pm sQ}(q_{\epsilon})=\left(-{{{\rm i}\over(sQ)^{2}\epsilon}}-{{1\over 2sQ}}+{{1\over(sQ)^{2}}}q{{\partial\over\partial q}}\right){{\hat{a}_{\pm sQ}(q)\over q^{\sigma_{\pm sQ}}}}+{\cal O}(\epsilon),$ (3.22a)
$a_{\pm sP}(q_{\epsilon,D})=\left({{{\rm i}\over(sQ)^{2}\epsilon}}+{{1\over sP}}\left({{{\rm i}\over 2\pi sQ}}-{1\over 2}\right)+{{1\over(sP)^{2}}}q_{D}{{\partial\over\partial q_{D}}}\right){{\hat{a}_{\pm sP}(q_{D})\over q_{D}^{\sigma_{\pm sP}}}}+{\cal O}(\epsilon),$ (3.22b)
respectively. We then require the cancellation of the $\epsilon$-poles in equations (3.22a) and (3.22b), which yields the relation
$\hat{a}_{\pm sQ}\bigl({\rm e}^{2\pi{\rm i}{P\over Q}}\bigr){\rm e}^{-2\pi{\rm i}{P\over Q}\sigma_{\pm sQ}}=\hat{a}_{\pm sP}\bigl({\rm e}^{2\pi{\rm i}{Q\over P}}\bigr){\rm e}^{-2\pi{\rm i}{Q\over P}\sigma_{\pm sP}},$ (3.23)
for all $s,P,Q\in{\mathbb{Z}}_{>0}$ with $P$, $Q$ coprime. We remark that, for the geometric branch, the formula in equation (3.23) assumes the simplified form [^10] A similar cancellation requirement was obtained in [44] in the study of the wave function for quantum mirror curves.
$\hat{a}^{({\rm geom})}_{sQ}\bigl({\rm e}^{2\pi{\rm i}{P\over Q}}\bigr)=\hat{a}^{({\rm geom})}_{sP}\bigl({\rm e}^{2\pi{\rm i}{Q\over P}}\bigr),$ (3.24)
@@
Let us now prove that the cancellation formula in equation (3.23) is a direct consequence of the integrality structure presented in equation (3.6). More precisely, equations (3.3) and (3.7) imply that
$\hat{a}_{n}(q)q^{-\sigma_{n}}=\sum_{k|n}kD_{k}\bigl(q^{n\over k}\bigr)q^{-{n\over k}\sigma_{k}}\sum_{j=0}^{k-1}q^{{n\over k}j},$ (3.25)
@@
$\hat{a}_{sQ}\bigl({\rm e}^{2\pi{\rm i}{P\over Q}}\bigr){\rm e}^{-2\pi{\rm i}{P\over Q}\sigma_{sQ}}=\sum_{k|sQ}kD_{k}\bigl({\rm e}^{2\pi{\rm i}{{sP\over k}}}\bigr){\rm e}^{-2\pi{\rm i}{{sP\over k}}\sigma_{k}}\sum_{j=0}^{k-1}{\rm e}^{2\pi{\rm i}{{sP\over k}}j}.$ (3.26)
Since the sum over $j$ on the right-hand side is non-zero if and only if $k|s$, equation (3.26) gives
$\hat{a}_{sQ}\bigl({\rm e}^{2\pi{\rm i}{P\over Q}}\bigr){\rm e}^{-2\pi{\rm i}{P\over Q}\sigma_{sQ}}=\sum_{k|s}k^{2}D_{k}(1)=c_{s}\in{\mathbb{Z}},\qquad s\in{\mathbb{Z}}_{>0},$ (3.27)
which is independent of the choice of $P,Q\in{\mathbb{Z}}_{>0}$ coprime, thus proving the cancellation identity in equation (3.23). We note that the case of $\hat{a}_{-n}(q)$ is handled analogously, yielding the same result. In particular, recalling that $\sigma_{-n}=0$, $n\in{\mathbb{Z}}_{>0}$, it defines the sequence of integers
$\hat{a}_{-sQ}\bigl({\rm e}^{2\pi{\rm i}{P\over Q}}\bigr)=\sum_{k|s}k^{2}D_{-k}(1)=c_{-s}\in{\mathbb{Z}},\qquad s\in{\mathbb{Z}}_{>0}.$ (3.28)
@@
$D_{\pm s}(1)={1\over s^{2}}\sum_{k|s}\mu(k)c_{\pm{s\over k}},\qquad s\in{\mathbb{Z}}_{>0},$ (3.29)
@@ 3.3 The wave function in the rational case, I
We will now apply the results obtained in Section 3.2 to derive explicitly the exact wave function $\chi(u;\hbar)$ in equation (3.13) as a formal power series in $X$, $X^{-1}$ for rational values of $\hbar$. Let us start from the holomorphic components $\phi^{(\alpha),{\rm WKB}}(X;q)$ and fix $\hbar$ as in equation (3.15), that is,
$q={\rm e}^{{\rm i}\hbar}={\rm e}^{2\pi{\rm i}{P\over Q}},\qquad P,Q\in{\mathbb{Z}}_{>0}\ {\rm coprime}.$ (3.30)
As before, we drop the label $\alpha$ for simplicity and intend the following discussion to hold for each non-abelian choice of the classical branch independently. The singular terms come from the values $n=sQ$ for $s\in{\mathbb{Z}}_{>0}$, while choosing $n=sQ+k$ with $1\leq k\leq Q-1$ produces only regular terms. Specifically, the terms involving singularities are obtained by substituting equation (3.22a) into equation (3.2) and contain the non-singular contributions
$\sum_{s\geq 1}\left(-{1\over 2}+{1\over sQ}q{\partial\over\partial q}\right){{\hat{a}_{\pm sQ}(q)\over sQq^{\sigma_{\pm sQ}}}}X^{\pm sQ}.$ (3.31)
@@
$\sum_{k=1}^{Q-1}{{1\over q^{k}-1}}\sum_{s\geq 0}{\hat{a}_{\pm(sQ+k)}(q)\over(sQ+k)q^{\sigma_{\pm(sQ+k)}}}X^{\pm(sQ+k)},$ (3.32)
where we have used that $q^{sQ+k}=q^{k}$. Thus, the sum of the quantities in equations (3.31) and (3.32) is the finite part of the function $\phi^{\rm WKB}(X;q)$ evaluated at the rational point in equation (3.15). It is convenient to introduce the functions
@@
$\varphi^{(+)}_{k}(X)$ $=\sum_{s\geq 0}{\hat{a}_{sQ+k}(q)\over(sQ+k)q^{\sigma_{sQ+k}}}X^{sQ+k},\qquad k=1,\dots,Q-1,$ (3.33c)
$\varphi^{(-)}_{k}(X)$ $=\sum_{s\geq 0}{\hat{a}_{-(sQ+k)}(q)\over-(sQ+k)q^{sQ+k+\sigma_{-(sQ+k)}}}X^{-(sQ+k)},\qquad k=1,\dots,Q-1,$ (3.33d)
@@
$\varphi_{k}(X)=\varphi^{(+)}_{k}(X)-q^{k}\varphi^{(-)}_{k}(X),\qquad k=1,\dots,Q-1.$ (3.34b)
Putting equations (3.31) and (3.32) together and suitably identifying the functions just introduced in equation (3.34), we can write the resummed series in the holomorphic block as
@@
Let us now consider the antiholomorphic component $\phi^{\rm WKB}(X_{D};q_{D})$ and fix $\hbar$ as above. The derivation, in this case, is very similar to the previous one except for one novelty, that is, the appearance of $\log X$ terms, which originate from the $\epsilon$-expansion of the $S$-dual variable
@@
Let us add a small positive term $\epsilon$ to the rational value of $\hbar$ as in equation (3.21). Then, we consider the limit $\epsilon\rightarrow 0$ and find that
$X_{\epsilon,D}=X^{2\pi\over\hbar_{\epsilon}}=X^{{Q\over P}{1\over 1+{\epsilon\over 2\pi}{Q\over P}}}=X^{Q\over P}\left(1-{Q^{2}\over 2\pi P^{2}}\epsilon\log X+{\cal O}\bigl(\epsilon^{2}\bigr)\right).$ (3.37)
As described before, the singular terms come from the values $n=sP$ for $s\in{\mathbb{Z}}_{>0}$, while choosing n=sP+k{n=sP+k} with $1\leq k\leq P-1$ produces only regular terms. Specifically, the terms involving singularities are obtained by substituting equations (3.22b) and (3.37) into equation (3.2) and contain the non-singular contributions
$\sum_{s\geq 1}\left({{{\rm i}\over 2\pi sQ}}-{1\over 2}+{1\over sP}q_{D}{\partial\over\partial q_{D}}\mp{{{\rm i}\over 2\pi}}\log X\right){{\hat{a}_{\pm sP}(q_{D})\over sPq_{D}^{\sigma_{\pm sP}}}}X^{\pm sQ}.$ (3.38)
@@
$\sum_{k=1}^{P-1}{{1\over q_{D}^{k}-1}}\sum_{s\geq 0}{\hat{a}_{\pm(sP+k)}(q_{D})\over(sP+k)q_{D}^{\sigma_{\pm(sP+k)}}}X^{\pm(sP+k){Q\over P}},$ (3.39)
where we have used that $q_{D}={\rm e}^{2\pi{Q\over P}}$ and thus $q_{D}^{sP+k}=q_{D}^{k}$. Therefore, the sum of the quantities in equations (3.38) and (3.39) is the finite part of the function $\phi^{\rm WKB}(X_{D};q_{D})$ evaluated at the rational point in equation (3.15). As we have done for the holomorphic part, we introduce the functions